A CI is desired for the true average stray-load loss μ (watts) for a certain type of induction motor when the line current is held at 10 amps for a speed of 1500 rpm. Assume that stray-load loss is normally distributed with σ = 3.3. (Round your answers to two decimal places.)

(a) Compute a 95% CI for μ when n = 25 and x = 57.5.

  ,

watts

(b) Compute a 95% CI for μ when n = 100 and x = 57.5.

  ,

watts

(c) Compute a 99% CI for μ when n = 100 and x = 57.5.

  ,

watts

(d) Compute an 82% CI for μ when n = 100 and x = 57.5.

  ,

watts

(e) How large must n be if the width of the 99% interval for μ is to be 1.0? (Round your answer up to the nearest whole number.)
n =
Answer #1

Solution :

Given that,

a) Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha/2 * ( \sigma /\sqrtn)

= 1.96 * ( 3.3 /  \sqrt25 )

= 1.29

At 95% confidence interval estimate of the population mean is,

\bar x  ± E

= 57.5  ± 1.29

= ( 56.21, 58.79 )

b) Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha/2 * ( \sigma /\sqrtn)

= 1.96 * ( 3.3 /  \sqrt100 )

= 0.65

At 95% confidence interval estimate of the population mean is,

\bar x  ± E

= 57.5  ± 0.65

= ( 56.85, 58.15 )

c) Z\alpha/2 = Z0.005 = 2.576

Margin of error = E = Z\alpha/2 * ( \sigma /\sqrtn)

= 2.576 * ( 3.3 /  \sqrt100 )

= 0.85

At 99% confidence interval estimate of the population mean is,

\bar x  ± E

= 57.5  ± 0.85

= ( 56.65, 58.35 )

d) Z\alpha/2 = Z0.09 = 1.341

Margin of error = E = Z\alpha/2 * ( \sigma /\sqrtn)

= 1.341 * ( 3.3 /  \sqrt100 )

= 0.44

At 82% confidence interval estimate of the population mean is,

\bar x  ± E

= 57.5  ± 0.44

= ( 57.06, 57.94 )

e) margin of error = E = width / 2 = 1.0 / 2 = 0.5

sample size = n = [Z\alpha/2* \sigma / E] 2

n = [ 2.576 * 3.3 / 0.5]2

n = 289.05

Sample size = n = 290